a streched string is observed to vibrate with frequency 30Hz in its fundamental mode .when the support are 60cm apart .the amplitude at the antinode is 3cm and the string has a mass of 30 kg . calculate the speed of propogation of the wave and the tension inthe string

Grade:

1 Answers

Ramesh Babu Gattu
12 years ago

dear friend,

this is for first part: answer: as it is fundamental mode, in stationary waves : L = wavelength/2 , so wavelength = 2L = 2*60cm = 120cm = 1.2m

and frequency = 30 Hz . so speed = frequency * wavelength = 30 * 1.2 = 36 m/s

this is the speed of propogation of the wave.

second part will be posted later.